how to find 5(2x-6)+2=2x+x
Found 2 solutions by checkley71, whatsmath?:
Answer by checkley71(8403)

5(2X-6)+2=2X+X
10X-30+2=3X
10X-3X=30-2
7X=28
X=28/7
X=4 ANSWER
PROOF
5(2*4-6)+2=2*4+4
5(8-6)+2=8+4
5*2+2=12
10=2=12
12=12

Answer by whatsmath?(5)

5(2X-6)+2=2X+X
10X-30+2=3X
10X-3X=30-2
7X=28
X=28/7
X=4